All past and current interventions at the Bioimaging Center
Dec 3, 2021, 2:56 PM
2-3.11.2021 - The WLL has been realigned and one of the internal HyDs exchanged.
Sep 16, 2021, 10:47 AM
16.09.2021 - The galvo-stage has been exchanged due to vibrations in it's motor. No more noise now.
Aug 18, 2021, 10:56 PM
23.08.2021 - There is some instability in the coupling of the WLL causing laser intensity loss. This parts will be exchanged on monday 23.
Aug 6, 2021, 9:40 AM
05.08.2021 : Service on the LSM780: The AOTF has been changed and all lasers have gained power (around 3X more and up to 10X for the 456nm, no change for the 405 which is unaffected by the AOTF).
Jun 17, 2021, 9:56 AM
This morning we realigned the laser fiber and could gain almost 3X output power ! We guess the fiber might have been inadvertently moved (at the laser box out-port, under the table). So you can expect now a much better signal from your samples!